# Efficient Hamiltonian, structure and trace distance learning of Gaussian states

### v2

@@ Abstract
In this work, we initiate the study of Hamiltonian learning for positive temperature bosonic Gaussian states, the quantum generalization of the widely studied problem of learning Gaussian graphical models. We obtain efficient protocols, both in sample and computational complexity, for the task of inferring the parameters of their underlying quadratic Hamiltonian under the assumption of bounded temperature, squeezing, displacement and maximal degree of the interaction graph. Our protocol only requires heterodyne measurements, which are often experimentally feasible, and has a sample complexity that scales logarithmically with the number of modes. Furthermore, we show that it is possible to learn the underlying interaction graph in a similar setting and sample complexity. Taken together, our results put the status of the quantum Hamiltonian learning problem for continuous variable systems in a much more advanced state when compared to spins, where state-of-the-art results are either unavailable or quantitatively inferior to ours. In addition, we use our techniques to obtain the first results on learning Gaussian states in trace distance with a quadratic scaling in precision and polynomial in the number of modes, albeit imposing certain restrictions on the Gaussian states. Our main technical innovations are several continuity bounds for the covariance and Hamiltonian matrix of a Gaussian state, which are of independent interest, combined with what we call the local inversion technique. In essence, the local inversion technique allows us to reliably infer the Hamiltonian of a Gaussian state by only estimating in parallel submatrices of the covariance matrix whose size scales with the desired precision, but not the number of modes. This way we bypass the need to obtain precise global estimates of the covariance matrix, controlling the sample complexity.
@@ 1 Introduction
Inferring parameters of quantum systems from measurement data is a fundamental task in quantum information science that has seen significant advancements over the recent years [HKT22a, AAKS20, BLMT24a, AA24, RF24, LTN+23, RSFOW24, SFMD+24, MBC+23, HTFS23, BLMT24b, Nar24, FGL+23, HKT+22b, LHT+24]. Although it has been known for roughly a decade that complete tomography of general quantum states inevitably requires a sample complexity scaling exponentially in the system’s size [FGLE12, OW16, HHJ+17], recent works have shown that it is possible to obtain a lot of information of practical interest much more efficiently. This is exemplified by the framework of classical shadows [HKP20].
Another fruitful approach to obtain efficient algorithms is to consider the learning of more restricted and physically motivated quantum states and evolutions, such as those generated by local Hamiltonians [LTN+23, SFMD+24, MBC+23, HTFS23, BLMT24b] or their Gibbs states [BLMT24a, HKT22a, AAKS20, RF24, Nar24, RSFOW24], where the problem of inferring the parameters of the underlying Hamiltonian is broadly referred to as the Hamiltonian learning problem. This line of work culminated in efficient algorithms (both in terms of samples and computationally) under various assumptions. The first sample-efficient algorithm for an arbitrary constant temperature local quantum Hamiltonian was given by [AAKS20], however the sample complexity was still polynomial in system size and the postprocessing was only efficient under further assumptions on the model, as it required solving a maximum entropy optimization problem. This was followed by work of [HKT22a], which focused on the high temperature regime. Resorting to cluster expansion techniques, they were able to obtain optimal sample complexity bounds and efficient post-processing. More recently [BLMT24a] gave Hamiltonian learning algorithms for general Hamiltonians with efficient sample and computational complexity based on sums-of-squares relaxations. However, the scaling of both the sample complexity and computational complexity is a high-degree polynomial and likely far from optimal.
@@
In the classical case there is also an extensive literature dedicated to the learning of Gaussian graphical models [Dem72, KKMM20, ZZ14, WRG16, FHT08, MVL20], which can be understood as continuous variable versions of discrete graphical models. In the last years, many fundamental questions related to this problem were settled, such as [MVL20], that gave sample optimal and efficient algorithms to learn such models, although the community still focuses on learning under more constrained settings, say in a differentially private way [AKT+23] and even much more complex models such as mixtures of Gaussians [BDJ+22]. The natural quantum variation of learning a Gaussian model is that of learning bosonic Gaussian states [Hol12], which correspond to Gibbs and ground states of quadratic Hamiltonians. Somewhat surprisingly given the prevalence and importance of such states in quantum optics and continuous variable approaches to quantum computing [Ser17], this problem was not considered prior to this work. Indeed, to the best of our knowledge, the only works that considered bosonic Hamiltonian learning problems focus on learning parameters from dynamics [LTN+23, MBC+23], not states.
@@ 2 Notations and Background
with $D=\oplus_{i=1}^{m}d_{i}I_{M_{2}}$, with $d_{i}>0$, and $S\in\mathrm{Sp}(2m,\mathbb{R})$. Let us also denote as $d_{\min}$ and $d_{\max}$ the smallest and the largest symplectic eigenvalues of $H$, respectively, and let $\lambda_{\min}$ and $\lambda_{\max}$ be the smallest and the largest eigenvalues of $H$, respectively. By Theorem 11 in [BJ15], $\lambda_{\min}\leq d_{\min}\leq d_{\max}\leq\lambda_{\max}$, and thus $1\leq\|S\|_{\infty}^{2}\leq\frac{\lambda_{\max}}{\lambda_{\min}}$. In the following, $\|S\|,d_{\min},d_{\max}$ will be used as parameters, as they have a natural physical interpretation, but these inequalities clarify how they relate to more usual parameterization of Hamiltonian matrices, in terms of their norm and condition number. We will say that a Hamiltonian matrix $H$ has graph $G=([m],\mathsf{E})$ made of $m$ vertices, where two vertices $i,j\in[m]$ are connected by an edge whenever the $2\times 2$ sub-matrix $\left\{H_{2i-\delta,2j-\delta^{\prime}}\right\}_{\delta,\delta^{\prime}\in\{0,1\}}$ is non-zero.
@@ 3 Main results
However, based on classical results [Dem72, KKMM20, ZZ14, WRG16, FHT08], we expect to be able to solve the Hamiltonian learning problem with a number of samples that scales logarithmically with the number of modes. To achieve such a scaling, we pioneer what we call the local inversion technique. The intuition behind the technique is that, for local Gibbs states, the correlations with neighboring sites should determine the interactions. For classical Gaussian graphical models this is true by conditional independence [Dem72, KKMM20, ZZ14, WRG16, FHT08], so it is possible to completely determine them by just considering constant-sized neighborhoods around each vertex. In contrast, we show that for Gaussian quantum states it suffices to post-process an estimate of $(2V-i\Omega)^{-1}$, constructed such that matrix entries corresponding to a pair of vertices are obtained from inverting submatrices of the covariance matrix corresponding to vertices that are at most $\sim\log(\epsilon^{-1})$ away from the vertices, to estimate the interactions with error up to $\epsilon$. To show this, we resort to various bounds on the sizes of blocks of inverses of block matrices and the Taylor series representation of $(2V-i\Omega)^{-1}$ in terms of the Hamiltonian. These bounds are explained in Sec. 6. But, crucially, the precision with which we need to estimate each of these submatrices to succesfully invert them locally does not scale with the number of modes, yielding our Hamiltonian learning result.
@@
$N=\mathcal{O}(\kappa^{-2-4\ln(\Delta)}\ln(m\delta^{-1}))$ (15)
@@ 3.1 Comments and future directions
First of all, we make a remark on the (implicit) dependence of our result on the condition number of $H$. In the sample complexity for Hamiltonian and graph learning, we see a dependence on $d_{\min}$, which can be interpreted as a “maximum temperature” parameter of the Gaussian state. Since $d_{\min}$ gives a smaller contribution to the Hamiltonian matrix, it is natural to ask if this dependence can be removed. In other quantum Hamiltonian learning results for discrete variable systems [BLMT24a, HKT22a, AAKS20, Nar24], even if with they display polynomial dependence on the system size, a dependence of this type does not appear. In fact, the dependence of our learning guarantees on the condition number can also be expressed in terms of potentially much smaller quantities, but we preferred to state our results as we did to avoid unnatural assumptions. Remarks 4.17.26.2 in the technical sections clarify this point. In the analogue classical problem of learning the precision matrix (the term used in this context for the Hamiltonian matrix) or the graph, the dependence on the condition number appeared in the first works on the subjects and was removed only by [MVL20], crucially using multiplicative error bounds on covariance estimation. In the graph learning problem, existence of edges is there formulated in terms of relative strength: two vertices $(i,j)$ are connected if Hamiltonian matrix element $H_{2i-\delta,2j-\delta^{\prime}}$ is sufficiently large compared with the diagonal elements $H_{2i-\delta,2i-\delta^{\prime}}$ and $H_{2j-\delta,2j-\delta^{\prime}}$. The sample complexity depends on the promise on the relative strength but not on the condition number. We leave exploring multiplicative error bounds in the quantum case for future work. See Appendix B for an argument suggesting that with additive precision guarantees one cannot avoid the condition number dependence.
@@ 4 Estimating covariance matrices and learning Gaussian states in trace distance
In this section we will discuss the sample and computational complexity of two fundnamental problems concerning the learning of Gaussian quantum states: estimating their covariance matrix and learning them in trace distance. The reason we will consider them together is that we show a new method to transfer a bound on the distance between covariance matrices to a bound in trace distance. Furthermore, estimates on the covariance matrix will be central to the other problems we consider in this paper, namely Hamiltonian and graph learning. Let us thus start with estimating the covariance matrix:
@@ 4.1 Estimating the entries of the covariance matrix
Arguing as before, we conclude that this SDP can also be solved in time that is polynomial in the number of modes and logarithmic in the desired precision $\epsilon$ and threhsold $\tau$. Again, this is feasilbe as $H$ itself is a solution. In addition, it is possible to only optimize over Hamiltonians that have the same underlying graph structure as $H$ by imposing the additional linear constraints $\widetilde{H}_{i,j}=0$ for $(i,j)\in\mathsf{E}$.
@@ 4.2 Learning in trace distance
Let $H=\frac{1}{2}\ln\left(\frac{2i\Omega V+I}{2i\Omega V-I}\right)i\Omega$ and $\hat{H}=\frac{1}{2}\ln\left(\frac{2i\Omega\hat{V}+I}{2i\Omega\hat{V}-I}\right)i\Omega$ for two bona fide covariance matrices $V$, $\hat{V}$. We also denote by $d_{\max}$, resp. $S$, the maximal symplectic eigenvalue of $H_{1}$, resp. the symplectic matrix associated to $H_{1}$ through Equation (67). Let $\rho$ and $\hat{\rho}$ be Gaussian states with $\rho\equiv\rho(t,H)$, $W\equiv\mathcal{W}(-\Omega^{-1}t)$, $\hat{\rho}\equiv\rho(\hat{t},\hat{H})$ and $\hat{W}\equiv\mathcal{W}(-\Omega^{-1}\hat{t})$. Given an approximation $\hat{V}$ of the covariance matrix $V$ as well as an approximation $\hat{t}$ of the mean vector $t$ to precision $\epsilon$: $\max_{i,j}|V_{i,j}-\hat{V}_{i,j}|\leq\epsilon$, and $\max_{i}|t_{i}-\hat{t}_{i}|\leq\epsilon$, such that $\|V-\hat{V}\|_{\infty}<\frac{e^{-2d_{\max}}}{1-e^{-2d_{\max}}}\|S\|_{\infty}^{-2}$ we have
@@
$\|H_{1}-H_{2}\|_{\infty}$ $\leq 4m\left(2\frac{e^{-2d_{\max}}}{1-e^{-2d_{\max}}}\|S\|_{\infty}^{-2}\right)^{-2}\epsilon=m(e^{2d_{\max}}-1)^{2}\|S\|_{\infty}^{4}\,.$
@@
Let $\rho(t,H)$ be a Gaussian state on $m$ modes s.t. the largest symplectic eigenvalue satisfies $d_{\max}$ satisfies $d_{\max}\leq\beta_{\max}$, the smallest symplectic eigenvalue $d_{\min}$ satisfies $d_{\min}\leq\beta_{\min}$, and $\|S\|_{\infty}\leq s$ for known $s,\beta_{\max},\beta_{\min}$. Then we have
@@ 5 Continuity bounds for Hamiltonians and covariances
Since $\|K_{1}-K_{2}\|_{\infty}<e^{-{2}d_{\max}}\sigma_{\min}(S^{\intercal}S)<\sigma_{\min}\left(\frac{t}{t+1}\frac{i\Omega V+I}{i\Omega V-I}+\frac{1}{t+1}\right)$ from (71), then

### v3

@@ Abstract
In this work, we initiate the study of Hamiltonian learning for positive temperature bosonic Gaussian states, the quantum generalization of the widely studied problem of learning Gaussian graphical models. We obtain efficient protocols, both in sample and computational complexity, for the task of inferring the parameters of their underlying quadratic Hamiltonian under the assumption of bounded temperature, squeezing, displacement and maximal degree of the interaction graph. Our protocol only requires heterodyne measurements, which are often experimentally feasible, and has a sample complexity that scales logarithmically with the number of modes. Furthermore, we show that it is possible to learn the underlying interaction graph in a similar setting and sample complexity. Taken together, our results put the status of the quantum Hamiltonian learning problem for continuous variable systems in a more advanced state when compared to spins, where state-of-the-art results are either unavailable or quantitatively inferior to ours. In addition, we use our techniques to obtain the first results on learning Gaussian states in trace distance with a quadratic scaling in precision and polynomial in the number of modes, albeit imposing certain restrictions on the Gaussian states. Our main technical innovations are several continuity bounds for the covariance and Hamiltonian matrix of a Gaussian state, which are of independent interest, combined with what we call the local inversion technique. In essence, the local inversion technique allows us to reliably infer the Hamiltonian of a Gaussian state by only estimating in parallel submatrices of the covariance matrix whose size scales with the desired precision, but not the number of modes. This way we bypass the need to obtain precise global estimates of the covariance matrix, controlling the sample complexity.
@@ 1 Introduction
Inferring parameters of quantum systems from measurement data is a fundamental task in quantum information science that has seen significant advancements over the recent years [HKT22a, AAKS20, BLMT24a, AA24, RF24, LTN+, RSFOW24, SFMD+24, MBC+23, HTFS23, BLMT24b, Nar24, FGL+23, HKT+22b, LHT+24]. Although it has been known for roughly a decade that complete tomography of general quantum states inevitably requires a sample complexity scaling exponentially in the system’s size [FGLE12, OW16, HHJ+17], recent works have shown that it is possible to obtain a lot of information of practical interest much more efficiently. This is exemplified by the framework of classical shadows [HKP20].
Another fruitful approach to obtain efficient algorithms is to consider the learning of more restricted and physically motivated quantum states and evolutions, such as those generated by local Hamiltonians [LTN+, SFMD+24, MBC+23, HTFS23, BLMT24b] or their Gibbs states [BLMT24a, HKT22a, AAKS20, RF24, Nar24, RSFOW24], where the problem of inferring the parameters of the underlying Hamiltonian is broadly referred to as the Hamiltonian learning problem. This line of work culminated in efficient algorithms (both in terms of samples and computationally) under various assumptions. The first sample-efficient algorithm for an arbitrary constant temperature local quantum Hamiltonian was given by [AAKS20], however the sample complexity was still polynomial in system size and the postprocessing was only efficient under further assumptions on the model, as it required solving a maximum entropy optimization problem. This was followed by work of [HKT22a], which focused on the high temperature regime. Resorting to cluster expansion techniques, they were able to obtain optimal sample complexity bounds and efficient post-processing. More recently [BLMT24a] gave Hamiltonian learning algorithms for general Hamiltonians with efficient sample and computational complexity based on sums-of-squares relaxations. However, the scaling of both the sample complexity and computational complexity is a high-degree polynomial and likely far from optimal.
@@
In the classical case there is also an extensive literature dedicated to the learning of Gaussian graphical models [Dem72, KKMM20, ZZ14, WRG16, FHT08, MVL20], which can be understood as continuous variable versions of discrete graphical models. In the last years, many fundamental questions related to this problem were settled, such as [MVL20], that gave sample optimal and efficient algorithms to learn such models, although the community still focuses on learning under more constrained settings, say in a differentially private way [AKT+23] and even much more complex models such as mixtures of Gaussians [BDJ+22]. The natural quantum variation of learning a Gaussian model is that of learning bosonic Gaussian states [Hol12], which correspond to Gibbs and ground states of quadratic Hamiltonians. Somewhat surprisingly given the prevalence and importance of such states in quantum optics and continuous variable approaches to quantum computing [Ser17], this problem was not considered prior to this work. Indeed, to the best of our knowledge, the only works that considered bosonic Hamiltonian learning problems focus on learning parameters from dynamics [LTN+, MBC+23], not states.
@@ 2 Notations and Background
with $D=\oplus_{i=1}^{m}d_{i}I_{M_{2}}$, with $d_{i}>0$, and $S\in\mathrm{Sp}(2m,\mathbb{R})$. Let us also denote as $d_{\min}$ and $d_{\max}$ the smallest and the largest symplectic eigenvalues of $H$, respectively, and let $\lambda_{\min}$ and $\lambda_{\max}$ be the smallest and the largest eigenvalues of $H$, respectively. By Theorem 11 in [BJ15], $\lambda_{\min}\leq d_{\min}\leq d_{\max}\leq\lambda_{\max}$, and since $\lambda_{\max}=\|H\|_{\infty}\geq\|S\|_{\infty}^{2}d_{\min}\geq\|S\|_{\infty}^{2}\lambda_{\min}$, we have $1\leq\|S\|_{\infty}^{2}\leq\frac{\lambda_{\max}}{\lambda_{\min}}$. In the following, $\|S\|,d_{\min},d_{\max}$ will be used as parameters, as they have a natural physical interpretation, but these inequalities clarify how they relate to more usual parameterization of Hamiltonian matrices, in terms of their norm and condition number. We will say that a Hamiltonian matrix $H$ has graph $G=([m],\mathsf{E})$ made of $m$ vertices, where two vertices $i,j\in[m]$ are connected by an edge whenever the $2\times 2$ sub-matrix $\left\{H_{2i-\delta,2j-\delta^{\prime}}\right\}_{\delta,\delta^{\prime}\in\{0,1\}}$ is non-zero.
@@ 3 Main results
However, based on classical results [Dem72, KKMM20, ZZ14, WRG16, FHT08], we expect to be able to solve the Hamiltonian learning problem with a number of samples that scales logarithmically with the number of modes. To achieve such a scaling, we pioneer what we call the local inversion technique. The intuition behind the technique is that, for local Gibbs states, the correlations with neighboring sites should determine the interactions. For classical Gaussian graphical models this is true by conditional independence [Dem72, KKMM20, ZZ14, WRG16, FHT08], so it is possible to completely determine them by just considering constant-sized neighborhoods around each vertex. In contrast, we show that for Gaussian quantum states it suffices to post-process an estimate of $(2V-i\Omega)^{-1}$, constructed such that matrix entries corresponding to a pair of vertices are obtained from inverting submatrices of the covariance matrix corresponding to vertices that are at most $\sim\log(\epsilon^{-1})$ away from the target pair of vertices, to estimate the interactions with error up to $\epsilon$. To show this, we resort to various bounds on the error in approximating inverses of approximately sparse matrices with local inversions, and the Taylor series representation of $(2V-i\Omega)^{-1}$ in terms of the Hamiltonian. These bounds are explained in Sec. 6. But, crucially, the precision with which we need to estimate each of these submatrices to succesfully invert them locally does not scale with the number of modes, yielding our Hamiltonian learning result.
@@
$N=\mathcal{O}(\kappa^{-2-6\ln(\Delta)}\ln(m\delta^{-1}))$ (15)
@@ 3.1 Comments and future directions
First of all, we make a remark on the (implicit) dependence of our result on the condition number of $H$. In the sample complexity for Hamiltonian and graph learning, we see a dependence on $d_{\min}$, which can be interpreted as a “maximum temperature” parameter of the Gaussian state. Since $d_{\min}$ gives a smaller contribution to the Hamiltonian matrix, it is natural to ask if this dependence can be removed. In other quantum Hamiltonian learning results for discrete variable systems [BLMT24a, HKT22a, AAKS20, Nar24], even if with they display polynomial dependence on the system size, a dependence of this type does not appear. In fact, the dependence of our learning guarantees on the condition number can also be expressed in terms of potentially much smaller quantities, but we preferred to state our results as we did to avoid unnatural assumptions. Remarks 4.17.26.2 in the technical sections clarify this point. In the analogue classical problem of learning the precision matrix (the term used in this context for the Hamiltonian matrix) or the graph, the dependence on the condition number appeared in the first works on the subjects and was removed only by [MVL20], crucially using multiplicative error bounds on covariance estimation. In the graph learning problem, existence of edges is there formulated in terms of relative strength: two vertices $(i,j)$ are connected if Hamiltonian matrix element $H_{2i-\delta,2j-\delta^{\prime}}$ is sufficiently large compared with the diagonal elements $H_{2i-\delta,2i-\delta^{\prime}}$ and $H_{2j-\delta,2j-\delta^{\prime}}$ (here $\delta,\,\delta^{\prime}$ can be 00 or $1$ and index position and momentum quadratures). The sample complexity depends on the promise on the relative strength but not on the condition number. We leave exploring multiplicative error bounds in the quantum case for future work. See Appendix B for an argument suggesting that with additive precision guarantees one cannot avoid the condition number dependence.
@@ 4 Estimating covariance matrices and learning Gaussian states in trace distance
In this section we will discuss the sample and computational complexity of two fundamental problems concerning the learning of Gaussian quantum states: estimating their covariance matrix and learning them in trace distance. The reason we will consider them together is that we show a new method to transfer a bound on the distance between covariance matrices to a bound in trace distance. Furthermore, estimates on the covariance matrix will be central to the other problems we consider in this paper, namely Hamiltonian and graph learning. Let us thus start with estimating the covariance matrix:
@@ 4.1 Estimating the entries of the covariance matrix
Arguing as before, we conclude that this SDP can also be solved in time that is polynomial in the number of modes and logarithmic in the desired precision $\epsilon$ and threhsold $\tau$. Again, this is feasible as $H$ itself is a solution. In addition, it is possible to only optimize over Hamiltonians that have the same underlying graph structure as $H$ by imposing the additional linear constraints $\widetilde{H}_{i,j}=0$ for $(i,j)\in\mathsf{E}$.
@@ 4.2 Learning in trace distance
Let $H=\frac{1}{2}\ln\left(\frac{2i\Omega V+I}{2i\Omega V-I}\right)i\Omega$ and $\hat{H}=\frac{1}{2}\ln\left(\frac{2i\Omega\hat{V}+I}{2i\Omega\hat{V}-I}\right)i\Omega$ for two bona fide covariance matrices $V$, $\hat{V}$. We also denote by $d_{\max}$, resp. $S$, the maximal symplectic eigenvalue of $H_{1}$, resp. the symplectic matrix associated to $H_{1}$ through Equation (68). Let $\rho$ and $\hat{\rho}$ be Gaussian states with $\rho\equiv\rho(t,H)$, $W\equiv\mathcal{W}(-\Omega^{-1}t)$, $\hat{\rho}\equiv\rho(\hat{t},\hat{H})$ and $\hat{W}\equiv\mathcal{W}(-\Omega^{-1}\hat{t})$. Given an approximation $\hat{V}$ of the covariance matrix $V$ as well as an approximation $\hat{t}$ of the mean vector $t$ to precision $\epsilon$: $\max_{i,j}|V_{i,j}-\hat{V}_{i,j}|\leq\epsilon$, and $\max_{i}|t_{i}-\hat{t}_{i}|\leq\epsilon$, such that $2\|V-\hat{V}\|_{\infty}<\frac{e^{-2d_{\max}}}{1-e^{-2d_{\max}}}\|S\|_{\infty}^{-2}$ we have
@@
$\|H-\hat{H}\|_{\infty}$ $\leq 2\left(2\frac{e^{-2d_{\max}}}{1-e^{-2d_{\max}}}\|S\|_{\infty}^{-2}\right)^{-1}\left(2\frac{e^{-2d_{\max}}}{1-e^{-2d_{\max}}}\|S\|_{\infty}^{-2}-2\|V-\hat{V}\|_{\infty}\right)^{-1}\|V-\hat{V}\|_{\infty}\,,$
@@
$\|H-\hat{H}\|_{\infty}$ $\leq 4m\left(2\frac{e^{-2d_{\max}}}{1-e^{-2d_{\max}}}\|S\|_{\infty}^{-2}\right)^{-2}\epsilon=m(e^{2d_{\max}}-1)^{2}\|S\|_{\infty}^{4}\epsilon\,.$
@@
Let $\rho(t,H)$ be a Gaussian state on $m$ modes s.t. the largest symplectic eigenvalue $d_{\max}$ satisfies $d_{\max}\leq\beta_{\max}$, the smallest symplectic eigenvalue $d_{\min}$ satisfies $d_{\min}\geq\beta_{\min}$, and $\|S\|_{\infty}\leq s$ for known $s,\beta_{\max},\beta_{\min}$. Then we have
@@ 5 Continuity bounds for Hamiltonians and covariances
Let $\sigma_{\min}(A)$ be the smallest singular vaule of $A$. Since $\|K_{1}-K_{2}\|_{\infty}<e^{-{2}d_{\max}}\sigma_{\min}(S^{\intercal}S)<\sigma_{\min}\left(\frac{t}{t+1}\frac{i\Omega V+I}{i\Omega V-I}+\frac{1}{t+1}\right)$ from (72), then